OpenStudy (anonymous):

If y varies directly with x and inversely with z, and y=16 when x=8 and z=2, find y when x=10 and z=5.

OpenStudy (saifoo.khan):

y = kx/z

OpenStudy (saifoo.khan):

16 = k8/2 k = 4

OpenStudy (anonymous):

Thanks!

OpenStudy (saifoo.khan):

y = (4)(10)/5 y = 8

OpenStudy (saifoo.khan):

Welks.